OpenStudy (anonymous):

Evaluate the limit. lim x approaches 0- x sqrt. (4 + x^−2)

OpenStudy (dumbcow):

\[x \sqrt{4+\frac{1}{x^{2}}} = x \sqrt{\frac{4x^{2} +1}{x^{2}}} = \sqrt{4x^{2} +1}\]

OpenStudy (anonymous):

@dumbcow What?

OpenStudy (dumbcow):

i simplified the expression so that limit could be evaluated at x=0

OpenStudy (dumbcow):

in its original form if you plug in 0 you get 0*infty which is indeterminate

OpenStudy (anonymous):

Good for you.

OpenStudy (anonymous):

so the answer is 1?

OpenStudy (anonymous):

but it's telling me it's not 1 :(

OpenStudy (anonymous):

\[\frac{x}{|x|} \sqrt{4x^2+1}\]

OpenStudy (dumbcow):

http://www.wolframalpha.com/input/?i=lim+x*sqrt%284%2Bx^-2%29+as+x-%3E0 ahh sorry i forgot to include neg since we are approaching from left

OpenStudy (anonymous):

but since it's factor shouldn't it be 1 regardless since it's x^2 muckushla: why the absolute bar?

OpenStudy (anonymous):

If you factor the 1/x^2 out rather than taking the x into the radical, you get the abs value in the denominator like @mukushla the result is that the limit from the left, does not equal the limit from the right and the limit does not exist... "dne"

OpenStudy (anonymous):

How to tell which way to go is the hard part. I see both ways working, but graphing on a calculator shows the dne result.

OpenStudy (anonymous):

I guess that taking the x into the radical loses information, whereas factoring out with the absolute value retains information.

OpenStudy (anonymous):

oh i see thank you